OpenStudy (anonymous):

Let f(x)=2x^-1, g(x)=2x+5 and H(x)=x-3. Perform the indicated operation. g(h(x))

OpenStudy (anonymous):

g(h(x)) = g(x-3) = 2 (x-3) +5 =2x-6+5 =2x-1